# A Diagnostic to Find and Help Combat Positivity Issues – with a Focus on Continuous Treatments

### v1

@@ Abstract
The positivity assumption is central in the identification of a causal effect, yet is rarely discussed, especially in conjunction with continuous treatments or Modified Treatment Policies. One common recommendation for dealing with a violation is to change the estimand. However, an applied researcher is faced with two problems: First, how can she tell whether there is a positivity violation given her estimand of interest, preferably without having to estimate a model first? Second, if she finds a problem with positivity, how should she change her estimand in order to arrive at an estimand which does not face the same issues? We suggest a novel diagnostic which allows the researcher to answer both questions by providing insights into how well an estimation for a certain estimand can be made for each observation using the data at hand. We provide a simulation study on the general behaviour of different MTPs at different levels of positivity violations and show how the diagnostic helps understand where bias is to be expected. We illustrate the application of our proposed diagnostic in a pharmacoepidemiological study based on data from CHAPAS-3, a trial comparing different treatment regimens for children living with HIV.
@@ 1 Introduction
Even though the assumption of positivity is essential to most standard causal estimation procedures, it is rarely discussed. In addition, most of the few works that consider consequences and solutions to positivity violations in detail are heavily focused on the binary treatment case, for example because “[…] it is the most commonly explored setting in the existing literature and has a well-developed set of methods available for comparison.” [1, page 1474]. However, continuous treatments are very often of interest in practice [2]. While positivity violations can be detected via cross tabulation in sufficiently simple binary settings, diagnosing positivity violations in a continuous setting is not trivial as there are always more values the variables can take than there are data points. This shows the need for a diagnostic that helps an applied researcher detecting positivity violations specifically in continuous data. Once a researcher has reached the conclusion that there are positivity issues with an estimand, a common recommendation is to switch the estimand [1]. Ideally, the researcher will find a compromise making sure the estimand is as close as possible to the original research question and avoiding bias due to positivity issues [3]. But this reveals another problem: How can a good alternative estimand be found, which solves this trade-off well?
@@
Regarding the identification of positivity issues, some tools exist which can be used in continuous treatment cases. Petersen et al. [3] suggested a parametric bootstrap which given an estimator produces an optimistic estimate of the bias. This tool may not always be very accurate and captures not only positivity issues but also other forms of sparsity, rendering its main purpose serving as a warning sign. The main drawback for this diagnostic is that it is estimator-specific. It can only be used once an estimator has been chosen and does not answer the question if the issue is with the estimator or the data. In addition, once positivity issues are identified with this tool, there is no guidance on what can be done to combat this issue. Another approach suggested by Zhu et al. [1] is to generalize propensity score based approaches, which are often used in order to deal with positivity issues in binary treatments [4], to continuous treatments. The generalized propensity score (GPS) is an estimate of the conditional density of the treatment given the variables of the adjustment set. However, this requires modeling a density first and leaves the researcher with a potentially high-dimensional density estimate, which cannot easily be used to derive information about positivity. In addition, the generalized propensity score assumes that observations with every feature combination require positivity in all potential treatments, but this is only true for a static treatment scheme. As discussed before, changing the estimand, typically to a more elaborate scheme, is one of the solutions to positivity problems, therefore a useful diagnostic needs to work on a large variety of estimands. Finally, the possibility of extrapolation from high density regions is not addressed at all if positivity is diagnosed via the GPS. Recently, Danelian et al. [5] proposed a regression tree based method to identify sparse areas called positivity regression tree (PoRT) algorithm. Chatton et al. [6] extended this to longitudinal settings with the sequential PoRT (sPoRT) algorithm. So far the authors have only considered binary exposure, but the algorithm could be adjusted to other types of treatment. However, as the algorithms depend on hyperparameters to define what constitutes a positivity issue, an extension to continuous treatments would mean this non-trivial question has to be answered first, likely falling back to ideas like the GPS.
@@
Once a researcher has diagnosed a positivity problem and whishes to change the estimand, they can use one of several estimands with relaxed positivity assumptions. For example, Kennedy [7] introduced a class of estimands called incremental propensity score interventions which do not rely on any positivity assumption as the propensity score itself is used to define the estimand. While this is potentially an interesting estimand to a researcher facing positivity issues, it solves the trade-off between bias due to positivity issues and closeness to the estimand of interest rather one-sidedly. In addition, the estimand is not necessarily easy to interpret and may not lead to a meaningful conclusion within the research field in question. Another large class of estimands with relaxed positivity assumptions are modified treatment policies (MTP) [8]. While MTP estimands leave the researcher a lot of options to define an estimand that can likely be estimated without suffering from a high bias due to positivity issues, it remains unclear, which MTPs will result in a satisfactorily low bias.
@@
In order to help applied researchers diagnose positivity issues and find good ways to respond to them, we introduce a diagnostic in this work. The diagnostic shows, given an estimand, how much support there is for all $n$ prediction problems of the causal estimand and helps a researcher compare different estimands based on which are likely to avoid bias due to positivity issues given a data set. The diagnostic can also act as a guide to a researcher aiming to tweak an estimand in order to avoid bias due to positivity issues.
@@
In the next chapter, the notation and setup of the work is introduced as well as core concepts. In Chapter 3, the diagnostic is first motivated and then two versions are introduced alongside a guide on how to choose hyperparameters. Chapter 4 contains a simulation study where the diagnostic is applied to different MTPs estimated on data with varying degrees of positivity violations, highlighting the issues a researcher faces when trying to choose an alternative estimand from this class. The diagnostic is applied to medical data in Chapter 5, where it is demonstrated how the diagnostic can be used to tweak problematic estimands. The paper concludes with a discussion of advantages and limitations in Chapter 6 and a conclusion in Chapter 7.
@@ 2.1 Notation and Setup
We consider an observed sample of $n$ i.i.d. observations of the form $O=(L,A,Y)\sim P_{0}$, where $L$ encompasses all adjustment variables needed to recover the causal quantity of interest, $A$ is the discrete or continuous treatment which we wish to intervene on, and $Y$ is the outcome variable. While the focus of this work lies on a standard outcome, the diagnostic which we introduce later on can easily be applied to, e. g., survival outcomes or multivariate $A$, as well. Following the potential outcomes framework [9], we denote the counterfactual outcome $Y$, which we would have observed had the treatment value been $a$, as $Y^{a}$. Instead of a fixed value $a$, corresponding to a static intervention, the treatment could also be determined by a function. For an overview of different types of intervention schemes see Section 2.3. In order to differentiate more easily between the observed sample and the intervened on sample, i. e., the sample with a changed value $a^{int}$ for $A$ according to the intervention scheme, we denote the observed values for a subject $i$ as $o_{i}^{obs}=(l_{i},a_{i}^{obs}:=a_{i})$ and the corresponding counterfactual values as $o_{i}^{int}=(l_{i},a_{i}^{int})$. The outcome $y_{i}$ is omitted here as it is not used in the diagnostic we present.
Given the three assumptions of consistency, positivity and unconfoundedness are fulfilled, there are multiple ways to recover a causal quantity of interest $\psi(P_{0})$; the focus of this work lies on three popular estimators, which we introduce here using the example of estimating the expected outcome of $Y$ under the intervention strategy [3]. Firstly, g-computation can be employed, which requires the estimate $\bar{Q}_{n}$, a model of $\mathbbm{E}(Y|A,L)$. Secondly, inverse probability of treatment weighting (IPTW) can be utilized, which requires an estimate $g_{n}$, a model of the probability of treatment $P_{0}(A=a|L)$ in the binary case or, respectively, $f_{0,A|L}(A|L)$ in the continuous case (the subscript zero will be dropped from the density subsequently). Lastly, doubly robust methods require the estimation of both. While other methods exist, they are either specialized to specific data structures (e. g., instrumental variable approaches) or rather basic (e. g., stratification), therefore we do not consider them in this work.
@@ 2.2.1 Definition
referring to a range of static interventions typically depicted in a causal dose-response curve [10]. In Section 2.3, this definition is extended to other intervention schemes. This positivity assumption postulates that, given possible covariate combinations, it is possible for an observation to receive all treatments of interest. The positivity assumption has been split into two different types: deterministic and stochastic positivity [1, 3]. Deterministic, or theoretical, violations encompass structural issues, like contraindication or illegality of a certain treatment. Stochastic, or practical, violations on the other hand cover issues, where for a certain subgroup of observations no or very little data are observed for some values of treatment. Therefore, a deterministic violation implies a stochastic violation. Independent of the type of positivity violation, this induces sparsity in the finite sample, which poses a challenge in estimation [3].
2.2.2 Responding to Positivity Violations
Current recommendations for dealing with sparsity due to positivity violations are extrapolation, if possible, and/or changing the estimand [1, 3]. The former requires using an estimator which allows for extrapolation [3]. The latter can be approached either by adjusting the target population (trimming/weighting the sample) [1, 3], or by adjusting the intervention scheme of interest, both of which represent a trade-off between proximity to the original research question of interest and identifiablity of the effect [3]. An alternative to extrapolation and changing the estimand is restricting $L$, resulting in a trade-off between bias due to positivity issues and bias due to confounding [3]. Within this work we will address all of these proposed solutions to stochastic positivity violations. The relation of positivity to inter- and extrapolation, which is a necessity in the continuous case, will be considered in detail in Section 3, where we introduce our diagnostic. The diagnostic also provides guidance on adjusting the target population, choosing a potential variable from the adjustment set to exclude from the analysis, and deciding between several potential estimands. However, in order to create alternative estimands, one has to be able to choose from a large variety of schemes in order to find the one providing the best trade-off between meeting your needs as a researcher and being estimable with the data at hand. Therefore, in the following we provide an overview of several types of intervention schemes, including Modified Treatment Policies or MTPs, which are a large class of intervention schemes from which several possible estimands can be created.
@@ 2.3 Advanced Treatment Schemes and Modified Treatment Policies
The simplest and most often used intervention scheme is the static intervention $A=a$, where $a$ is the same fixed value for all subjects in the study. Sometimes many values for $a$ are of interest and the goal is to estimate a Causal Dose Response Curve (CDRC), where for each value of $a$ the counterfactual outcome $Y^{a}$ is presented in a graph [2]. For this type of intervention the strict positivity assumption as stated above needs to be met. The class of static interventions can be extended to dynamic interventions, where the value of the treatment depends on variables of the adjustment set $l$ leading to $A=\mathbbm{d}(l)$, for some function $\mathbbm{d}$ [8]. The positivity assumption for a dynamic scheme is less strict compared to a fixed scheme. For example, if a certain treatment value $a$ only occurs for a sub-group of the sample with $L=l$, $f(a|L)$ only needs to be positive for $L=l$. Modified Treatment Policies (MTP) extend dynamic schemes by letting the function $\mathbbm{d}$ depend on the natural value of $A$, leading to $A=\mathbbm{d}(l,a^{obs})$ [8]. This relaxes the positivity assumption further. Doubly robust estimation of MTP causal estimands as described by Díaz et al. [8] can be conducted using the R-Package lmtp [11].
@@
where $E$ are combinations of $a^{int},l$ which are seen in the estimand (see Appendix A). For example in a dynamic scheme, where women are given a dose of 5 g of medication and men receive 3 g and no other variables exist, positivity is only needed for $f(a=5|\text{woman})$ and $f(a=3|\text{man})$. Therefore MTPs relax the positivity assumption by giving the researcher more freedom to define a treatment scheme which avoids sparse data regions and represent an important tool in the toolkit of every researcher trying to combat positivity issues.
@@ 3 A Kernel-based Sparsity Diagnostic
In this section, we introduce a feasibility diagnostic, which provides guidance to applied researchers about whether a causal estimand can, in principle, be estimated given a specific data set. The diagnostic is designed to be estimand-specific, because the positivity assumption depends on the estimand, and data set-specific, because practical positivity violations are a property of the data, not necessarily the data generating process. The diagnostic is also designed to be estimator-independent, since as long as there is one estimator that can recover the causal quantity of interest, the research question can be answered. In order to construct this diagnostic, the sources of bias related to positivity issues in estimation are examined in the following.
@@ 3.1 Sources of Bias
The first source of bias is often addressed by using doubly robust estimation with a large variety of highly adaptive models in the super learner [12]. This is the only source of bias named by Petersen et al. which is not linked to positivity and is therefore not addressed further. The second source of bias represents the deterministic positivity violation and means there are areas in the feature space where it is impossible to observe data. This is, as will be discussed in the next subsection, rare and can be mitigated using inter- and extrapolation from neighboring areas where data was observed, albeit by introducing another assumption, for example in the form of a parametric model. Inter- and extrapolation are also a necessity when dealing with the third form of bias, finite sample bias for $g_{n}$ and $\bar{Q}_{n}$, at least in a continuous setting, as a finite sample means there is no data almost everywhere. Finite sample bias is an issue in any finite data set, but the problem is worsened in sparse regions of the data. Therefore, practical positivity violations lead to an increase in finite sample bias. In a continuous setting, both 2 and 3 force us to rely on inter- and extrapolation, the only relevant difference in estimation between the two is that 3 could theoretically be mitigated by collecting more data. For a given data set, however, they can be treated as the same problem: sparsity [3]. The last source of bias stems from low estimates of $g_{n}$ or, respectively, truncation of $g_{n}$ in response to low estimates amplifying any existing bias. This last bias is estimator-specific, as it is only relevant for models which are using $g_{n}$. However, this includes all state-of-the-art doubly-robust methods. Doubly-robust methods are effectively reduced to g-computation when the the last source of bias is too high [3]. But, while this means one cannot reap the benefits of this model class, if this is the only problematic source of bias, it is not relevant for estimability as $\bar{Q}_{n}$ is consistent and does not suffer from sparsity issues. However, as both sparsity and consistency are not black-and-white categories, it is to be expected that this last source of bias will slightly increase bias for a doubly-robust estimator even if consistency and sparsity for the estimation of $\bar{Q}_{n}$ are satisfactory, compared to a world where the estimation for both $\bar{Q}_{n}$ and $g_{n}$ is satisfactory in this regard.
@@
As the first and last source of bias are estimator-specific, meaning they can in theory be completely avoided by choosing a different estimator, the second and third source, which can be summarized as sparsity, are targeted in our diagnostic. The last source of bias is always targeted indirectly, but is addressed more explicitly in a second more complex variation of the diagnostic. It is important to note that not all of the 3 biases are caused by (practical) violations of positivity. But since all types of finite sample bias threaten the estimability of our estimand given the data set in the same way as finite sample bias due to positivity issues, i. e., due to the lack of data where it is needed, considering positivity violations together with other types of sparsity in our feasibility diagnostic is preferable to limiting this diagnostic to strictly positivity-induced issues.
3.2 Extrapolation and Positivity

### v3

# A Diagnostic to Find and Help Combat Stochastic Positivity Issues – with a Focus on Continuous Treatments

@@ Abstract
The positivity assumption is central in the identification of a causal effect. Especially its stochastic variant is an issue many applied researchers face. Yet positivity is rarely discussed, especially in conjunction with continuous treatments or Modified Treatment Policies. One common recommendation for dealing with a violation is to change the estimand. However, an applied researcher is faced with two problems: First, how can she tell whether there is a stochastic positivity violation given her estimand of interest, preferably without having to estimate a model first? Second, if she finds a problem with stochastic positivity, how should she change her estimand in order to arrive at an estimand which does not face the same issues? We suggest a novel diagnostic which allows the researcher to answer both questions by providing insights into how well an estimation for a certain estimand can be made for each observation using the data at hand. We provide a simulation study on the general behaviour of different Modified Treatment Policies (MTPs) at different levels of stochastic positivity violations and show how the diagnostic helps understand where bias is to be expected. We illustrate the application of our proposed diagnostic in a pharmacoepidemiological study based on data from CHAPAS-3, a trial comparing different treatment regimens for children living with HIV.
@@ 1 Introduction
Even though the assumption of positivity is essential to most standard causal effect estimation procedures, it is rarely discussed. In addition, most of the few works that consider consequences and solutions to positivity violations in detail are heavily focused on the binary treatment case, for example because “[…] it is the most commonly explored setting in the existing literature and has a well-developed set of methods available for comparison.” [1, page 1474]. However, continuous treatments are very often of interest in practice [2]. Theoretical positivity violations can often be identified using logic, as one of the typical examples, an intervention where the treatment hysterectomy is assigned to a man without a womb [3], illustrates. Stochastic positivity violations on the other hand can be detected via cross tabulation in sufficiently simple binary settings. But diagnosing them in a continuous setting is not trivial as there are always more values the variables can take than there are data points. This shows the need for a diagnostic that helps an applied researcher detect stochastic positivity violations specifically in continuous and high-dimensional data. Once a researcher has reached the conclusion that there are stochastic positivity issues with an estimand, a common recommendation is to switch the estimand [1]. Ideally, the researcher will find a compromise making sure the estimand is as close as possible to the original research question and avoiding bias due to positivity issues [4]. But this reveals another problem: How can a good alternative estimand, which solves this trade-off well, be found?
@@
Regarding the identification of stochastic positivity issues, some tools exist which can be used in continuous treatment cases. Petersen et al. [4] suggested a parametric bootstrap which, given an estimator, produces an optimistic estimate of the bias. This tool may not always be very accurate and captures not only positivity issues but also other forms of sparsity, rendering its main purpose serving as a warning sign. The main drawback for this diagnostic is that it is estimator-specific. It can only be used once an estimator has been chosen and does not answer the question if the issue is with the estimator or the data. In addition, once stochastic positivity issues are identified with this tool, there is no guidance on what can be done to combat this issue. Another approach suggested by Zhu et al. [1] is to generalize propensity score based approaches, which are often used in order to deal with stochastic positivity issues in binary treatments [5], to continuous treatments. The generalized propensity score (GPS) is an estimate of the conditional density of the treatment given the variables of the adjustment set. However, this requires modeling a density first and leaves the researcher with a potentially high-dimensional density estimate, which cannot easily be used to derive information about positivity. In addition, the generalized propensity score assumes that observations with every feature combination require positivity in all potential treatments, but this is only true for a static treatment scheme. As discussed before, changing the estimand, typically to a more elaborate scheme, is one of the solutions to positivity problems, therefore a useful diagnostic needs to work on a large variety of estimands. Finally, the possibility of extrapolation from high density regions is not addressed at all if stochastic positivity is diagnosed via the GPS. Recently, Danelian et al. [6] proposed a regression tree based method to identify sparse areas called positivity regression tree (PoRT) algorithm. Chatton et al. [7] extended this to longitudinal settings with the sequential PoRT (sPoRT) algorithm. So far the authors have only considered binary exposure, but the algorithm could be adjusted to other types of treatment. However, as the algorithms depend on hyperparameters to define what constitutes a stochastic positivity issue, an extension to continuous treatments would mean this non-trivial question has to be answered first, likely falling back to ideas like the GPS and some threshold value.
@@
Once a researcher has identified a stochastic positivity problem and wishes to change the estimand, they can use one of several estimands with relaxed positivity assumptions. For example, Kennedy [8] introduced a class of estimands called incremental propensity score interventions which do not rely on any positivity assumption as the propensity score itself is used to define the estimand. While this is potentially an interesting estimand to a researcher facing positivity issues, it solves the trade-off between bias due to stochastic positivity issues and closeness to the estimand of interest rather one-sidedly. In addition, the estimand is not necessarily easy to interpret and may not lead to a meaningful conclusion within the research field in question. Another large class of estimands with relaxed positivity assumptions are modified treatment policies (MTP) [9]. While MTP estimands leave the researcher a lot of options to define an estimand that can likely be estimated without suffering from a high bias due to stochastic positivity issues, it remains unclear which MTPs will result in a satisfactorily low bias.
@@
In order to help applied researchers diagnose stochastic positivity issues and find effective ways to respond to them, we introduce a diagnostic in this work. The diagnostic shows, given an estimand, how much support there is for all $n$ prediction problems of the causal estimand and helps a researcher compare different estimands based on which are likely to avoid bias due to stochastic positivity issues given a data set. The diagnostic can also act as a guide to a researcher aiming to tweak an estimand in order to avoid bias due to stochastic positivity issues.
@@
In the next section, the notation and setup of the work is introduced as well as core concepts. In Section 3, the diagnostic is first motivated and then two versions are introduced alongside a guide on how to choose hyperparameters. Section 4 contains a simulation study where the diagnostic is applied to different MTPs estimated on data with varying degrees of stochastic positivity violations, highlighting the issues a researcher faces when trying to choose an alternative estimand from this class. The diagnostic is applied to medical data in Section 5, where it is demonstrated how the diagnostic can be used to tweak problematic estimands. The paper concludes with a discussion of advantages and limitations in Section 6 and a conclusion in Section 7.
@@ 2.1 Notation and Setup
We consider an observed sample of $n$ independent and identically distributed (i.i.d.) observations of the form $O=(L,A,Y)\sim P_{0}$, where $L$ encompasses all adjustment variables needed to recover the causal quantity of interest, $A$ is the discrete or continuous treatment which we wish to intervene on, and $Y$ is the outcome variable. While the focus of this work lies on a standard outcome and univariate treatment, the diagnostic, which we introduce later on, can easily be applied to, e. g., survival outcomes, or multivariate $A$ as well. Following the potential outcomes framework [10], we denote the counterfactual outcome $Y$, which we would have observed had the treatment value been $a$, as $Y^{a}$. Instead of a fixed value $a$, corresponding to a static intervention, the treatment could also be determined by a function. For an overview of different types of intervention schemes see Section 2.3. In order to differentiate more easily between the observed sample and the intervened on sample, i. e., the sample with a changed value $a^{int}$ for $A$ according to the intervention scheme, we denote the observed values for a subject $i$ as $o_{i}^{obs}=(l_{i},a_{i}^{obs})$ and the corresponding counterfactual values as $o_{i}^{int}=(l_{i},a_{i}^{int})$. The outcome $y_{i}$ is omitted here as it is not used in the diagnostic we present.
Given the three assumptions of consistency, (theoretical) positivity and unconfoundedness are fulfilled, there are multiple ways to recover a causal quantity of interest $\psi(P_{0})$ [4]. The focus of this work lies on three popular estimators, which we introduce here using the example of estimating the expected outcome of $Y$ under the intervention strategy. Firstly, g-computation can be employed, which requires the estimate $\bar{Q}_{n}$, a model of $\mathbbm{E}(Y|A,L)$. Secondly, inverse probability of treatment weighting (IPTW) can be utilized, which requires an estimate $g_{n}$, a model of the probability of treatment $P_{0}(A=a|L)$ in the binary case or, respectively, $f_{0,A|L}(A|L)$ in the continuous case (the subscript zero will be dropped from the density subsequently for better readability). Lastly, doubly robust methods require the estimation of both. While other methods exist, they are less general, typically either specialized to specific data structures (e. g., instrumental variable approaches [11]), rather basic (e. g., stratification), or take a completely different point of view on estimation, like some Bayesian approaches [12], therefore we do not consider them in this work.
@@ 2.2.1 Definition
referring to a range of static interventions typically depicted in a causal dose-response curve [13]. In Section 2.3, this definition is extended to other intervention schemes. This positivity assumption postulates that, given possible covariate combinations, it is possible for an observation to receive all treatments of interest. The positivity assumption has been split into two different types: deterministic and stochastic positivity [1, 4]. Deterministic, or theoretical, violations encompass especially illogical estimands, for example, a pregnancy in a person without a female reproductive system, and are therefore not an estimation issue per se. Stochastic, or practical, violations on the other hand cover issues, where for a certain subgroup of observations no or very little data are observed for some values of treatment, either due to low probabilities in the data generating process or simply chance. Stochastic positivity violations result in sparsity in the finite sample, which poses a challenge in estimation as they go hand in hand with a high bias [4]. Note that we use the term sparsity in this work not as a descriptor for a problem stemming from high-dimensionality of the data alone, but as an issue related to local sample size, which is influenced by the sample size, the dimensionality of the data, and the distribution of the data.
2.2.2 Responding to Stochastic Positivity Violations
Current recommendations for dealing with sparsity due to stochastic positivity violations are extrapolation (in fact, if perfect extrapolation can be assumed, the assumption of positivity is not needed anymore [14]), if possible, and/or changing the estimand [1, 4]. The former requires using an estimator which allows for extrapolation [4]. The latter can be approached either by adjusting the target population (trimming/weighting the sample) [1, 4], or by adjusting the intervention scheme of interest, both of which represent a trade-off between proximity to the original research question of interest and identifiability of the effect [4]. An alternative to extrapolation and changing the estimand is restricting $L$, resulting in a trade-off between bias due to stochastic positivity issues and bias due to confounding [4]. Within this work we will address all of these proposed solutions to stochastic positivity violations. The relation of stochastic positivity to inter- and extrapolation, which is a necessity in the continuous case, will be considered in detail in Section 3, where we introduce our diagnostic. The diagnostic provides guidance on adjusting the target population, choosing a potential variable from the adjustment set to exclude from the analysis, and deciding between several potential estimands. However, in order to create alternative estimands, one has to be able to choose from a large variety of schemes in order to find the one providing the best trade-off between meeting your needs as a researcher and being estimable with the data at hand. Therefore, in the following we provide an overview of several types of intervention schemes, including Modified Treatment Policies (MTPs), which are a large class of intervention schemes from which several estimands can be created.
@@ 2.3 Advanced Treatment Schemes and Modified Treatment Policies
The simplest and most often used intervention scheme is the static intervention $A=a$, where $a$ is the same fixed value for all subjects in the study. Sometimes many values for $a$ are of interest and the goal is to estimate a Causal Dose Response Curve (CDRC), where for each value of $a$ the counterfactual outcome $Y^{a}$ is presented in a graph [2]. For this type of intervention the strict positivity assumption as stated above needs to be met. The class of static interventions can be extended to dynamic interventions, where the value of the treatment depends on variables of the adjustment set $l$ leading to $A=\mathbbm{d}(l)$, for some function $\mathbbm{d}$ [9]. The positivity assumption for a dynamic scheme is less strict compared to a fixed scheme. For example, if a certain treatment value $a$ only occurs for a sub-group of the sample with $L=l$ in the estimand, $f(a|L)$ only needs to be positive for $L=l$. Modified Treatment Policies (MTP) extend dynamic schemes by letting the function $\mathbbm{d}$ depend on the natural value of $A$, leading to $A=\mathbbm{d}(l,a^{obs})$ [9]. This relaxes the positivity assumption further. Doubly robust estimation of MTP causal estimands as described by Díaz et al. [9] can be conducted using the R-Package lmtp [15].
@@
where $E$ are combinations of $a^{int}$ and $l$ which are seen in the estimand (see Appendix A). For example in a dynamic scheme, where women are given a dose of 5 g of medication and men receive 3 g and no other variables exist, positivity is only needed for $f(a=5|\text{woman})$ and $f(a=3|\text{man})$. Therefore MTPs relax the positivity assumption by giving the researcher more freedom to define a treatment scheme which avoids sparse data regions and represents an important tool in the toolkit of every researcher trying to combat stochastic positivity issues.
@@ 3 A Kernel-based Sparsity Diagnostic
In this section, we introduce a feasibility diagnostic, which provides guidance to applied researchers about whether a causal estimand can, in principle, be estimated given a specific data set. The diagnostic is designed to be estimand-specific, because the positivity assumption depends on the estimand, and data set-specific, because stochastic positivity violations are a property of the data, not necessarily the data generating process. The diagnostic is also designed to be estimator-independent, since as long as there is one estimator that can recover the causal quantity of interest, the research question can be answered. In order to construct this diagnostic, the sources of bias related to positivity issues in estimation are examined in the following.
@@ 3.1 Sources of Bias
The first source of bias is often addressed by using doubly robust estimation with a large variety of highly adaptive models in the super learner [16]. This is the only source of bias named by Petersen et al. which is not linked to positivity and is therefore not addressed further. The second source of bias represents the deterministic positivity violation and means there are areas in the feature space where it is impossible to observe data. This can be mitigated using inter- and extrapolation from neighboring areas where data was observed, albeit by introducing another assumption, for example in the form of a parametric model. Inter- and extrapolation are also a necessity when dealing with the third form of bias, finite sample bias for $g_{n}$ and $\bar{Q}_{n}$, at least in a continuous setting, as a finite sample means there is no data almost everywhere. Finite sample bias is an issue in any finite data set, but the problem is exacerbated in sparse regions of the data. Therefore, practical positivity violations lead to an increase in finite sample bias. In a continuous setting, both 2 and 3 force us to rely on inter- and extrapolation, the only relevant difference in estimation between the two is that 3 could theoretically be mitigated by collecting more data. For a given data set, however, they can be treated as the same problem: sparsity [4]. The last source of bias stems from low estimates of $g_{n}$ or, respectively, truncation of $g_{n}$ in response to low estimates amplifying any existing bias. This last bias is estimator-specific, as it is only relevant for estimators which are using $g_{n}$. However, this includes all state-of-the-art doubly robust methods. Doubly robust methods are effectively reduced to g-computation when the last source of bias is too high [4]. But, while this means one cannot reap the benefits of this model class, if this is the only problematic source of bias, it is not relevant for estimability as $\bar{Q}_{n}$ is consistent and does not suffer from sparsity issues. However, as both sparsity and consistency are not black-and-white categories in practice, it is to be expected that this last source of bias will slightly increase bias for a doubly robust estimator. This holds, even if consistency and sparsity for the estimation of $\bar{Q}_{n}$ are satisfactory, compared to a world where the estimation for both $\bar{Q}_{n}$ and $g_{n}$ is satisfactory in this regard.
@@
As the first and last source of bias are estimator-specific, meaning they can in theory be completely avoided by choosing a different estimator, the second and third source, which can be summarized as sparsity, are targeted in our diagnostic. The last source of bias is always targeted indirectly, but is addressed more explicitly in a second more complex variation of the diagnostic. It is important to note that not all of the 3 biases are caused by stochastic violations of positivity alone. But since all types of finite sample bias threaten the estimability of our estimand given the data set in the same way as finite sample bias due to stochastic positivity issues, i. e., due to the lack of data where it is needed, considering stochastic positivity violations together with other types of sparsity in our feasibility diagnostic is preferable to limiting this diagnostic to strictly positivity-induced issues.
3.2 Extrapolation and Stochastic Positivity